OpenStudy (anonymous):

Use a half angle identity to find the exact value of the expression: sin105 degrees

OpenStudy (anonymous):

well, sin(105) = sin(45+55) = (sin addition formula). But then we would still need the exact value of sin55. But we can play the same trick using 55 = 45 + 10.

OpenStudy (anonymous):

I have the formula sin u/2= +- radical 1-cos u/ 2 is that correct?

OpenStudy (anonymous):

We can make it sin (60+45)

OpenStudy (anonymous):

But how do I plug it into that formula? Or is that the correct formula?

OpenStudy (anonymous):

What formula? do you have specific formula?

OpenStudy (anonymous):

@amandamv1994

OpenStudy (anonymous):

It says to use half angle identity. But in my book it says half angle formula. Its sin u/2 =+- radical 1-cos u/2

OpenStudy (anonymous):

I don't think you can use this formula here because you have to break the angle to known angles which are 0,30,45,60,90,180,270 or 360

OpenStudy (anonymous):

Hm okay. Cause the answer is supposed to be \[\sqrt{2-\sqrt{3}}\div 2\]

OpenStudy (anonymous):

do you have formula for sin(60+45) and how can you break it? Did you take these stuff?

OpenStudy (anonymous):

Is that the sin (u+v) ?

OpenStudy (anonymous):

Addition and Subtraction Formulas For Sine and Cosine 1. sin(x + y) = sin(x)cos(y) + cos(x)sin(y) 2. sin(x – y) = sin(x)cos(y) − cos(x)sin(y) 3. cos(x + y) = cos(x)cos(y) − sin(x)sin(y) 4. cos(x – y) = cos(x)cos(y) + sin(x)sin(y) Did you take those equations before?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

we can apply the first one in your case here Which is : sin(x + y) = sin(x)cos(y) + cos(x)sin(y)

OpenStudy (anonymous):

sin(60+ 45) = sin(60) cos(45) + cos(60) sin(45)

OpenStudy (anonymous):

Then thats when we refer to the unit circle?

OpenStudy (anonymous):

\[(\sqrt{3}/2)*(1/\sqrt{2}) + (0.5 * (1/\sqrt{2})\]

OpenStudy (anonymous):

I guess so I don't really remember the name but let's see if we get the same answer

OpenStudy (anonymous):

Cause the points for 45 degrees is \[\sqrt{2/2, \sqrt{2/2}}\]

OpenStudy (anonymous):